# Share Your Secrets for Privacy! Confidential Forecasting with Vertical Federated Learning

### v1

@@ 3.1 Protocol Overview
An overview of STV is provided in Algorithm 1. The framework consists of preliminary steps for pre-processing the series and secretly sharing the input features (Figure 2), followed by decentralized training and inference (Figure 3).
Training. To start, the active party initiates by pre-processing the output and determines parameters like the auto-correlations and partial auto-correlations of the series (line 2). As illustrated in 2(a), these are then used to generate a polynomial for SARIMAX or ARTs (line 3). As illustrated in 2(b), the features and outputs are then secretly shared and transformed into lagged design matrices (line 7), like in Equation 2. All parties then follow decentralized training protocols, Algorithm 2 or Algorithm 3 to train a distributed model. Training is illustrated in 3(a).
Privacy-preserving inference. During inference (3(b)), the final prediction exists as a distributed share (line 14), which is aggregated on the requesting party (line 16). Since the result is computed as a distributed share, the outputs are not produced by a sole party, which makes the inference serverless.
@@ 3.2 STVT: Autoregressive tree (ART)
STVT focuses on the autoregressive tree and the first step is to transform the original datasets, $X$ and $Y$, into time-lagged design matrices. Following the training framework in Xie et al. [42], we train a distributed autoregressive XGBoost model using the secretly-shared design matrices. The details are provided in Algorithm 2.
Training proceeds iteratively, finally resulting in the generation of $T$ trees on each party. At each step, every party learns a new tree and makes a local prediction (lines 4-5). The active participant aggregates the distributed shares of the prediction to compute the first and second order gradients needed for generating the next distributed tree. These gradients are then secretly shared with the other parties for them update their local tree ensembles. Details on the tree-building function, SecureFit, are provided in Appendix C.
During training, individual predictions are aggregated on the active party since gradients are computed by $C_{1}$ (see Appendix C). For inference, aggregation can be performed on any party since gradients are not calculated. This enables serverless forecasting.
@@ 3.3 STVL: SARIMAX
With STVL, the objective is optimizing the coefficients, $A$, in Equation 2. This can be done either directly or iteratively using the two-step regression process explained in Section 2.1. Details are provided in Algorithm 3.
Several steps in Algorithm 3 require matrix operations like multiplications (lines 8-9; 13-19) and inverse (line 14) on secretly shared data. This is a requirement for both iterative and direct methods, as seen in Equation 5 and Equation 4. We design novel $N$-party ($\geq 2$) algorithms for performing these operations, as detailed below.
N-party matrix multiplication. To perform matrix multiplication with secretly shared data (algorithm 4), we view the computation of every output element $W_{i,j}$ as a scalar product of row and column vectors (line 3). This can be implemented using the $N$-party element-wise product of the row and column vectors with Beaver’s triples followed by an addition operation, as shown in Section 2.2.
N-party matrix inverse. For matrix inverses (Algorithm 5), we compute the inverse of a secretly-shared matrix, $U$, using a non-singular perturbation matrix, $P$, generated by the active party (line 2). Subsequently, the aggregation of product $UP$ on a passive party does not leak $U$ as the party does not know $P$ (line 5). $(UP)^{-1}$ can then be computed locally and secretly shared, followed by a matrix multiplication with $P$, i.e., $P\times(UP)^{-1}=U^{-1}$, giving the result as a secret share (lines 7-10).
Similar to the tree-based models, forecasting is done by secretly sharing features and computing the output as a distributed share (see 3(b)). The true value of the forecast can be obtained by summing the shares across all parties, enabling serverless inference.
4 Experiments
We evaluate the forecasting accuracy of STV against centralized approaches. We also compare the scalability of iterative and direct optimization for linear forecasters using the total communication cost.
@@ 4 Experiments
Figure 4: Forecasts from one of the windows of total size 50 (40 train, 10 test) from three datasets.
@@ 4.1 Forecasting accuracy
We compare the performance of STVL (direct optimization) and STVT with other centralized methods: Long-Short-Term Memory (LSTM), SARIMAX with MLE [^1] https://www.statsmodels.org/devel/generated/statsmodels.tsa
@@
, and diffusion models for forecasting (SSSDS4[2].
The LSTM has two layers of 64 LSTM units each, followed by two dense layers of size 32 and 1. We train the model for 500 epochs and use a batch size of 32 with a default learning rate of 0.001. The diffusion configuration has the following settings [^2] Using https://github.com/AI4HealthUOL/SSSD.git: $T=200,\beta_{0}=0.0001,\beta_{T}=0.02$. For the diffusion model, we use two residual layers, four residual and skip channels, with three diffusion embedding layers of dimensions $8\times 16$. Training is done for 4000 iterations with a learning rate of 0.002.
Five public forecasting datasets are used: Airline passengers [20], Air quality data [39], PV Power [17], SML 2010 [30], and Rossman Sales [1]. An industry-specific dataset to estimate a performance parameter for chip overlays from inline sensor values in semiconductor manufacturing is also included [11, 12]. Additional details on the datasets, evaluation, and pre-processing are given in Appendix A and Appendix B.
A variant of prequential window testing [5] is used since industrial time series data can significantly change after intervals due to machine changes/repairs. The dataset is partitioned into multiple windows of a given size, each further divided into an 80-20 train-test split. After forecasting the test split, the model is retrained on the next window (refer Appendix B).
All features and outputs are scaled between 0 and 1 for a consistent comparison. We thus present the predictions’ normalized mean-squared errors ($n$-MSE). Ground truths on the test set are measured and averaged across multiple windows. We average the $n$-MSE scores across different window sizes to generalize performance scores across varying forecasting ranges, as shown in Table 2.
Table 2 demonstrates that our approach STVL outperforms other centralized methods in terms of performance. We use direct optimization since iterative gradient descent eventually converges to the same value in the long run. Due to its guaranteed convergence, direct optimization improves against centralized methods by up to 23.81%. Though it does worse than SARIMAX MLE on the passenger data, it is still among the top two methods.
We also show three regression plots from a prequential window of size 50 from three datasets in Figure 4. In general, all the methods can capture patterns in the time series, like in 4(a) and 4(b). However, deep models such as SSSDS4 occasionally overfit, showing the drawback of NNs on small windows (see 4(c)).
@@ 4.2 Scalability
We measure the total communication costs of direct optimization using the normal equation (NE) and iterative batch gradient descent (GD) to analyze the scalability of the two methods under increasing parties, features, and samples.
We vary the parties between 2, 4, and 8, features between 10 and 100, and samples between 10, 100, and 1000. Since communication costs depend only on the dataset dimensions and the number of parties, we generate random data matrices for all valid combinations of features and samples on each party, i.e., #features $\leq$ #samples. We then optimize the coefficients using either the direct approach (Algorithm 3, lines (13-19)), or batch gradient descent (lines (6-10)), for a different number of iterations (10, 100, 1000). For party-scaling, we average the total communication costs across various feature and sample combinations for a given number of parties. Similarly, we measure feature and sample scaling by averaging the total costs across different (sample, party) and (feature, party) combinations. Results are shown in Table 3, Table 4, and Table 5.
Table 3: Average total communication sizes (bytes) under varying party count.
parties NE GD iterations
@@
2 2.54×$\times 10^{8}$ 2.33×$\times 10^{7}$ 2.33×$\times 10^{8}$ 2.33×$\times 10^{9}$
@@
Table 4: Average total communication sizes (bytes) with varying feature counts.
Features NE GD iterations
10 100 1000
10 9.77×$\times 10^{6}$ 8.34×$\times 10^{6}$ 8.34×$\times 10^{7}$ 8.34×$\times 10^{8}$
@@
Table 5: Average of total communication sizes (bytes) with varying samples.
Samples NE GD iterations
10 100 1000
10 1.16×$\times 10^{6}$ 2.76×$\times 10^{5}$ 2.76×$\times 10^{6}$ 2.76×$\times 10^{7}$
@@
Our observation from these three tables is that when the number of parties/samples/features is small, direct optimization’s cost is comparable to an iterative version with more iterations. For example, when the number of parties is 2, we see that the cost of NE is close to GD with 100 iterations in Table 3. However, when we increase to 8 parties, the cost of NE exceeds GD with 100 iterations.
Similarly, in Table 4 and Table 5, we see that NE has a lower cost than GD with 100 iterations for smaller magnitudes of features and samples. This is no longer the case when increasing the features and samples.
@@
However, the cost of GD increases proportionally with the iterations, and at some point, it becomes more expensive than NE, as we see with 1000 iterations. In practice, hyperparameters like the learning rate affect the steps required for convergence, which may be hard to tune in a distributed setup. So, choosing between iterative or direct optimization depends on several factors, requiring adaptability in frameworks.
5 Related Work
We direct our attention to the selected works presented in Table 6 to represent closely related research in VFL comprehensively. We compare the methods’ applicability to time series forecasting, whether inference privacy can be easily achieved through decentralization, their adaptability to alternative optimization choices, and whether they can generalize to $N$-parties.
Time-series forecasting. Earlier, we mentioned that industries require easy-to-understand and low-complexity models to avoid overfitting on small datasets. In this light, we focus on linear/logistic regression (LR) [16, 14, 34], and tree-based models [42, 8], with transparent structures that are considered inherently interpretable [31, 45]. Yan et al. [43] modifies the split learning architecture using Gated Recurrent Units (GRUs) with a shared upper model for predictions. However, its dependence on a single server for generating predictions is a bottleneck and can lead to trust issues among the involved parties. Nevertheless, to the best of our knowledge, this remains the only other method for time-series forecasting with VFL.
Serverless inference. We consider schemes adopting SS as potential candidates for serverless inference requirements, as these can be seamlessly integrated with a decentralized approach akin to ours [42, 34, 14]. Shi et al. [34] and Han et al. [14] train linear models while Xie et al. [42] implement XGBoost. Homomorphic encryption methods, such as Hardy et al. [16], require the predictions to be decrypted at the server, violating the requirement.
Optimization. When it comes to optimization techniques, all selected works, except for Han et al. [14], employ solely iterative approaches. Notably, Han et al. [14] offer iterative and matrix-based methods for direct optimization using Equation 5. This requires computing matrix multiplications and inverses, only implemented for the two-party case. To the best of our knowledge, ours is the first $N$-party protocol for performing these operations.
Table 6: Comparison of related works in VFL.
Method
Time Series
Serverless inference
Dual optimization
N-party ($\geq 2$)
  Yan et al. [43]
Han et al.[14]
Xie et al.[42]
Shi et al.[34]
STV (ours)
6 Applications
In this work, Additionally, the challenge of slow data collection is also prevalent in domains as often real datasets need to be collected via manually conducting surveys or interviews.
7 Conclusion and Discussion
This work presents a novel decentralized and privacy-preserving forecasting framework. Here, we focus on the manufacturing industry as a motivating example for our problem scenario. However, our framework’s applicability also extends to other domains, such as healthcare and finance, which face similar constraints. For example, collaborating healthcare centers can jointly benefit from sharing patient features to model patient risk levels. However, patient confidentiality agreements may pose restrictions to sharing data.
Our results show that our approach is competitive against centralized methods. Scalability analyses show the nuanced dynamics between iterative and direct optimization, highlighting the need for an adaptable framework.
For the future, several paths of exploration beckon. First, the models used in this study are weaker at capturing long-range patterns in time series data, for which LSTMs and diffusion models are better suited. We aim to enable these models for VFL. Second, current VFL cases assume a static set of participants. Exploring a hybrid of horizontal FL and VFL elements would be interesting if VFL model clusters were independently trained and aggregated to enhance their collective forecasting power. Going beyond semi-honest participants to malicious adversaries is another possibility.

### v2

@@ 4.1 Protocol Overview
An overview of STV is provided in Algorithm 1. The framework consists of preliminary steps for pre-processing the series and secretly sharing features (Figure 2), followed by distributed training and inference (Figure 3).
Training. The active party initiates by pre-processing the output and determines parameters like the auto-correlations and partial auto-correlations of the series (line 2) to identify the time lags of the autoregressive and moving average terms in eq.(1). As shown in 2(a), these are then used to generate a polynomial for SARIMAX or ARTs (line 3). As illustrated in 2(b), the features and outputs are then secretly shared and transformed into lagged design matrices (line 7), like in eq.(2). All parties then follow decentralized training protocols, Algorithm 2 or Algorithm 3 to train a distributed model. Training is illustrated in 3(a).
Privacy-preserving inference. During inference (3(b)), the final prediction exists as a distributed share (line 14), which is aggregated on the requesting party (line 16). Since the result is computed as a distributed share, the outputs are not tied to a particular party.
@@ 4.2 STVT: Autoregressive tree (ART)
STVT focuses on the autoregressive tree and transforms the original datasets, $X$ and $Y$, into time-lagged design matrices. Following the training framework in Xie et al. [44], we train a distributed autoregressive XGBoost model using the secretly-shared design matrices. The details are provided in Algorithm 2. Training proceeds iteratively, finally resulting in the generation of $T$ trees on each party. At each step, every party learns a new tree and makes a local prediction (lines 4-5). The active participant aggregates the distributed shares of the prediction to compute the first and second order gradients needed for generating the next distributed tree. These gradients are then secretly shared with the other parties for them update their local tree ensembles. Details on the tree-building function, SecureFit, are provided in Appendix C. During training, individual predictions are aggregated on the active party since gradients are computed by $C_{1}$ (see Appendix C). For inference, aggregation can be performed on any party since gradients are not calculated.
@@ 4.3 STVL: SARIMAX
With STVL, the objective is optimizing the coefficients, $A$, in Equation 2. This can be done either analytically (exactly) or iteratively using the two-step regression process explained in Section 2.1. Details are provided in Algorithm 3.
Algorithm 3 requires matrix operations like multiplication (lines 8-9; 13-19) and inverse (line 14) on secretly shared data. This is a requirement for both iterative and exact methods, as seen in Equation 4 and Equation 5. The $N$-party ($\geq 2$) algorithms for performing these operations are detailed below.
N-party matrix multiplication. To perform matrix multiplication with secretly shared data (algorithm 4), we view the computation of every output element $W_{i,j}$ as a scalar product of row and column vectors (line 3). This is implemented using the $N$-party element-wise product of the row and column vectors with Beaver’s triples followed by an addition, as shown in Section 2.2.
N-party matrix inverse. As shown in Algorithm 5, we compute the inverse of a secretly-shared matrix, $U$, using a non-singular perturbation matrix, $P$, generated by the active party (line 2). Subsequently, the aggregation of product $UP$ on a passive party does not leak $U$ as $P$ is unknown (line 5). $(UP)^{-1}$ can then be computed locally and secretly shared, followed by a matrix multiplication with $P$, i.e., $P\times(UP)^{-1}=U^{-1}$, giving the result as a secret share (lines 7-10).
Similar to the tree-based models, forecasting is done by secretly sharing features and computing the output as a distributed share (see 3(b)). The true value of the forecast can be obtained by summing the shares across all parties, avoiding the need for having a server act as a middleman for producing forecasts.
5 Performance Evaluation
We evaluate the forecasting accuracy of STV against centralized approaches. We also compare the scalability of iterative and exact optimization for linear forecasters using the total communication cost.
@@ 5 Performance Evaluation
Figure 4: Forecasts (y-axis) of different methods, total window size 50 (40 steps horizon and forecasts 10 steps ahead). The x-axis shows the forecasted timesteps.
@@ 5.1 Forecasting accuracy
We compare the performance of STVL (exact optimization) and STVT with other centralized methods: Long-Short-Term Memory (LSTM), SARIMAX with MLE [^1] https://www.statsmodels.org/devel/generated/statsmodels.tsa
@@
, and diffusion models for forecasting (SSSDS4[2]. The LSTM has two layers with 64 units, followed by two dense layers of sizes 32 and 1. We train the model for 500 epochs with a batch size of 32 and a learning rate of 0.001. The diffusion configuration has the following settings [^2] Using https://github.com/AI4HealthUOL/SSSD.git: $T=200,\beta_{0}=0.0001,\beta_{T}=0.02$. The model uses two residual layers, four residual and skip channels, with three diffusion embedding layers of dimensions $8\times 16$. Training is done for 4000 iterations with a learning rate of 0.002.
Five public forecasting datasets are used: Airline passengers [21], Air quality data [43], PV Power [18], SML 2010 [34], and Rossman Sales [1]. An industry-specific dataset to estimate a performance parameter for chip overlays from inline sensor values in semiconductor manufacturing is also included [13, 12]. Evaluation and pre-processing steps are given in Appendix A and Appendix B. A variant of prequential window testing [5] is used since industrial time series data can significantly change after intervals due to machine changes/repairs. The dataset is partitioned into multiple windows of a given size, each further divided into an 80-20 train-test split. After forecasting the test split, the model is retrained on the next window (cf. Appendix B). All features and outputs are scaled between 0 and 1 for a consistent comparison. We thus present the predictions’ normalized mean-squared errors ($n$-MSE). Ground truths on the test set are measured and averaged across multiple windows. We average the $n$-MSE scores across different window sizes to generalize performance scores across varying forecasting ranges, as shown in Table 3.
Table 3 demonstrates that our approach STVL outperforms other centralized methods in terms of performance. We use exact optimization since iterative gradient descent eventually converges to the same value in the long run. Due to its guaranteed convergence, exact optimization improves against centralized methods by up to 23.81%. Though it does worse than SARIMAX MLE on the passenger data, it is still among the top two methods. Regression plots from a prequential window of size 50 from three datasets are shown in Figure 4. Generally, all methods can capture patterns in the time series, like in 4(a) and 4(b). Occasionally, deep models such as SSSDS4 overfit, showing the drawback of NNs on small windows (see 4(c)).
@@ 5.2 Scalability
We measure the total communication costs of exact optimization using the normal equation (NE) and iterative batch gradient descent (GD) to analyze the scalability of the two methods under increasing parties, features, and samples. We vary the parties between 2, 4, and 8, features between 10 and 100, and samples between 10, 100, and 1000. Since communication costs depend only on the dataset dimensions and the number of parties, we generate random data matrices for all valid combinations of features and samples on each party, i.e., #features $\leq$ #samples. We then optimize the coefficients using either the exact approach (Algorithm 3, lines (13-19)), or batch gradient descent (lines (6-10)), for a different number of iterations (10, 100, 1000). For party-scaling, we average the total communication costs across various feature and sample combinations for a given number of parties. Similarly, we measure feature and sample scaling by averaging the total costs across different (sample, party) and (feature, party) combinations. Results are shown in Table 4.
Table 4: Network overhead (B) with varying parties, feature counts and samples.
NE GD iterations
@@
Parties 2 2.54×$\times 10^{8}$ 2.33×$\times 10^{7}$ 2.33×$\times 10^{8}$ 2.33×$\times 10^{9}$
@@
Features 10 9.77×$\times 10^{6}$ 8.34×$\times 10^{6}$ 8.34×$\times 10^{7}$ 8.34×$\times 10^{8}$
@@
Samples 10 1.16×$\times 10^{6}$ 2.76×$\times 10^{5}$ 2.76×$\times 10^{6}$ 2.76×$\times 10^{7}$
@@
We observe that when the number of parties/samples/features is small, exact optimization’s cost is comparable to an iterative version with more iterations. For example, when the number of parties is 2, we see that the cost of NE is close to GD with 100 iterations. However, when we increase to 8 parties, the cost of NE exceeds GD with 100 iterations. Similarly, we see that NE has a lower cost than GD with 100 iterations for smaller magnitudes of features and samples. This is no longer the case when increasing the features and samples. However, the cost of GD increases proportionally with the iterations, eventually surpassing NE, as seen with 1000 iterations. In practice, hyperparameters like the learning rate and batch size affect the convergence rate, which may be hard to tune in a distributed setup. So, choosing between iterative or exact optimization depends on several factors, requiring adaptability in frameworks.
6 Conclusion
We present a generalizable, novel decentralized and privacy-preserving forecasting framework. For instance, healthcare centers can share patient data to assess risks, but confidentiality agreements limit data sharing. Our results show that the approach competes well with centralized methods. Scalability analyses reveal the trade-offs between iterative and exact optimization, emphasizing the need for adaptability. Future directions include enabling deep models, such as LSTMs and diffusion, in VFL and exploring hybrid approaches that combine horizontal and vertical FL for improved forecasting. Additionally, moving beyond semi-honest to adversarial participants is a potential area of exploration.
{credits}
Acknowledgements
This research is part of the Priv-GSyn project, 200021E_229204 of Swiss National Science Foundation and the DEPMAT project, P20-22 / N21022, of the research programme Perspectief which is partly financed by the Dutch Research Council (NWO).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Datasets and Pre-processing
We make use of the following public datasets in our work: Air Quality [43], SML 2010 [34], PV Power [18], Airline Passengers [21], and Rossman Sales [1]. We do not provide the details on the industrial dataset due to confidentiality agreements. Here we provide a brief description of each dataset and the pre-processing steps applied to the dataset itself. These steps are unconnected with the series pre-processing steps which are part of the STV framework. For all datasets, we scale all features and output values between the range 0 to 1 using a MinMax scaler to ensure that all datasets have the same range of values for comparing the MSE losses.
0.A.1 Air Quality
The dataset contains approximately 9300 samples of multivariate time-series data, with 15 attributes. Five of these are true output values on five gases: Carbon Monoxide (CO), Non Metanic Hydrocarbons (NMHC)), Benzene (C6H6)), Total Nitrogen Oxides (NOx), and Nitrogen Dioxide (NO2). Exogenous features such as the temperature, ozone levels, and humidity are provided, along with strongly correlated sensor data for each of the five gases. The dataset contains missing values and duplicates, and contains hourly data for each of the five gases.
We preprocess the data by discarding all rows with any missing information and remove duplicate rows. We predict the ground truth values of CO using the other sensor values and information such as temperature and humidity as the exogenous regressors.
0.A.2 SML 2010
The SML 2010 dataset contains infomation from a monitor system in a domotic house. It contains approximately 4100 samples with 24 attributes in total, corresponding to 40 days of monitoring data. The attributes contain values such as the indoor and outdoor temperature, lighting levels, Carbon Dioxide levels, relative humidity, rain, windspeed, etc. We predict the indoor habitation temperature using the others as exogenous features.
0.A.3 Airline Passengers
Airline Passengers is a small dataset of 145 samples containing the number of international airline passengers (in thousands) on a monthly basis. The exogenous features are also just two: the year and the month. We predict the number of passengers using the year and month as exogenous regressors.
0.A.4 PV Power
The PV Power dataset contains around 3100 samples of solar power generation data from each of two power plants over a 34-day period. Attributes include features such as the DC power, AC power, yield, ambient temperature, irradiation levels, and the data and time.
We drop identifiers, empty, and duplicate data. We also drop the DC power attribute, and total yield as these features are very strongly correlated with the AC output. As outputs, we predict the AC power generation using the remaining features as exogenous regressors.
0.A.5 Rossman Sales
The Rossman Sales dataset contains sales data for 1115 store outlets. The attributes consists of features such as holidays, store type, competitor distance, number of customers, and promotional details among others. We predict the sales of the store with ID 1, using the other features as exogenous regressors.
Appendix 0.B Experimental Evaluation
As mentioned in the main text, we use a variation of prequential window testing [5], whereby the entire data is broken into windows of a defined length, each one internally split in an 80-20 train-test ratio. This is illustrated in Figure 5.
Figure 5: Prequential window evaluation with re-training in every window
For each window, we train on the portion allotted for training and forecast the remaining. Within a given training window, we first generate the polynomial by processing the time series as in Algorithm 1 from the main text. Identifying the parameters for generating the polynomial can be automated using implementations such as auto arima [^3] https://alkaline-ml.com/pmdarima/modules/generated/pmdarima.
arima.auto_arima.html
For each window size, such as 50, 100, 200, 400, we average the MSE loss between the forecasts and the true values across all windows. The average MSE per-window size is given in Table 5, which is an expanded version of Table 1 from the main text.
  Dataset Window size STVL SARIMAX MLE LSTM STVT $SSSD^{S4}$
Air quality 50 0.00071 0.00110 0.00244 0.00198 0.00270
100 0.00059 0.00048 0.00055 0.00075 0.00086
200 0.00066 0.00124 0.00069 0.00067 0.00100
400 0.00080 0.00068 0.00087 0.00061 0.00144
Avg. 0.00069 0.00088 0.00114 0.00100 0.00150
SML 2010 50 0.00645 0.00621 0.00343 0.00755 0.01373
100 0.01958 0.02819 0.02778 0.04305 0.02384
200 0.00500 0.00662 0.02781 0.01709 0.00336
400 0.00045 0.00030 0.00960 0.00729 0.00249
Avg. 0.00787 0.01033 0.01716 0.01875 0.01085
Rossman Sales 50 0.00070 0.00079 0.00464 0.00183 0.00589
100 0.00064 0.00071 0.00340 0.00496 0.00281
200 0.00068 0.00086 0.00674 0.00088 0.00233
400 0.00095 0.00072 0.01079 0.00206 0.00220
Avg. 0.00074 0.00077 0.00639 0.00243 0.00331
PV Power 25 0.00133 0.00093 0.00950 0.00660 0.00131
50 0.00360 0.00307 0.01037 0.00258 0.00143
100 0.00004 0.00063 0.01115 0.00010 0.00267
200 0.00054 0.00175 0.54227 0.00068 0.00128
Avg. 0.00138 0.00159 0.14333 0.00249 0.00167
Airline Passengers 60 0.00261 0.00113 0.11651 0.00673 0.00110
80 0.00450 0.00444 0.00339 0.00628 0.00403
100 0.00308 0.00066 0.01982 0.00270 0.00798
120 0.00316 0.00136 0.05164 0.01134 0.00356
140 0.00185 0.00351 0.01512 0.01332 0.00293
Avg. 0.00304 0.00222 0.04130 0.00808 0.00392
Table 5: Average normalized MSE values for different public datasets, with different prequential window sizes
Appendix 0.C SMPC-based XGBoost with VFL
XGBoost [6] is a tree-based gradient-boosting algorithm, that iteratively generates an ensemble of trees by greedily learning a new tree at every step to improve on the earlier one. Each tree has weights assigned to its leaf nodes. When making a prediction for a sample, the weights corresponding to the leaf to which the sample was assigned to are summed to give the final prediction score.
To generate a tree at every iteration, it uses first and second order gradients of the latest predictions, i.e., from the previous tree, in order to set the optimal weights for the new one.
For each sample with index $i$, the first and second order gradients are denoted as $g_{i}$ and $h_{i}$, respectively. The sum of the gradients of all instances on a particular node is used to set the new weights for it. For example, for node $j$ and corresponding instance set $I_{j}$, the accumulated values of $g$ and $h$ are computed as follows: $G_{j}=\sum_{i\in I_{j}}g_{i}$, and $H_{j}=\sum_{i\in I_{j}}h_{i}$. Based on this, for a tree with $T$ nodes, the weights and objective are calculated as follows:
$w_{j}=-G_{j}/(H_{j}+\lambda)$ (7)
$obj=-0.5\times\sum_{j=1}^{T}((G_{j})^{2}/(H_{j}+\lambda))+\gamma T$ (8)
, where $\gamma,\lambda$ are regularizers. While eq.(7) sets the new weights, eq.(8) is used to identify how to split nodes at each iteration.
With this in mind, using the secret sharing primitives it is possible to compute these functions to extend XGBoost to VFL, which we show in Algorithm 6.
When XGBoost is trained for VFL using secret sharing, each client obtains a local tree with their own weights as shown in Figure 6. In the figure, the weights for clients 1 and 2 are distributed such that their local weights are shares of the weights of the centralized version, i.e., $wi=w{i1}+w{i2}$ $\forall i\in[1,4]$
The indicator vector $s$ on line 1 of algorithm 6, is a binary vector that is used to point out the location of instances on nodes. We explain this with the help of the example in Figure 7. To calculate the updated weight of the node with instances that have an age greater than 30 (bottom left), we need to find the sum of $\sum_{i\in I_{j}}g_{i}$, where $I_{j}=\{2,4\}$ eq.(7). The indicator vector in this case would be $s=[0,1,0,1]$, meaning that nodes 2 and 4 are part of node $j$. If we have a vector of the gradients, $g=[g1,g2,g3,g4]$, we can compute $g_{2}+g_{4}$ as $s\odot g$, i.e., the inner product. Under VFL, both the gradients $g$, and the indicator vector $s$, exist as secret shares across clients, i.e., $s=\sum_{k}^{K}$$\langle s_{k}\rangle^{k}$, and $g=\sum_{k}^{K}$$\langle g_{k}\rangle^{k}$. Therefore, we can compute the product using MPC primitives for matrix multiplication.
Figure 6: Centralized vs Secretly shared XGBoost trees
Figure 7: Mapped instances to a node in XGBoost
The process of split-finding and setting weights is done using the SecureBuild function, which makes use of secret sharing primitives to compute the functions in eq.(8) and eq.(7). We defer readers to the implementation in Xie et al. [44] for additional details on this.
Algorithm 6 Fit XGBoost [44]
Data: Secretly shared matrices: $\langle X\rangle$, $\langle Y\rangle$, $\langle\hat{Y}\rangle$ across $K$ clients, $C_{1},C_{2},..,C_{K}$
Output: Learned tree for the current iteration, one on each client.
1:  Initialize indicator vector $s\leftarrow 1$ on all $C_{k}$
2:  if active client $C_{1}$ then
3:     Compute derivatives $g,h$
4:     Generate shares $\langle g\rangle$, $\langle h\rangle$, $\langle s\rangle$
5:  end if
6:  $Tree_{k}=\textbf{\emph{SecureBuild}}$($\langle g\rangle^{k}$,$\langle h\rangle^{k}$, $\langle s\rangle^{k}$) on each $C_{k}$
7:  return  $Tree_{k}$ on $C_{k}$